Whoa I need help please?
The equation y=-1777x+27,153 can be used to predict the number y of gun death in the U.S. x years after 2000, that is x=0 corresponds to 2000, x=2 corresponds to 2002, x=4 corresponds to 2004 and so on, predict the number of gun deaths in 2004 and 2007. In what year will the number of gun deaths be 11,160?
2004 ?
2007 ?
and the number of gun deaths will be 11,160 in the year ?
I am totally lost. Thank you

Answer by Theo(13342)

equation is:

y = -1777*x + 27153

x = 0 corresponds to the year 2000.

you want to predict the number of gun deaths in 2004 and 2007.

2004 corresponds to x = 4
2007 corresponds to x = 7

in 2004, the equation becomes y = -1777*4 + 27153 = 20045
in 2007, the equation becomes y = -1777*7 + 27153 = 14714

to solve for what year the total number of gun deaths will be 11160, you need to substitute for y in the equation to get:

11160 = -1777*x + 27153
subtract 27153 from both sides of this equation to get:
11160 - 27153 = -1777*x
combine like terms to get:
-15993 = -1777*x
divide both sides by -1777 to get:
-15993/-1777 = x
solve for x to get:
x = 9

the year 2000 + 9 = the year 2009

your answers are:

number of gun deaths in 2004 = 20045
number of gun deaths in 2007 = 14714
gun deaths will be 11160 in 2009